# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ 5 Improving Label DP Training with Retraining (RT)
Motivated by our theoretical results in Section 4 which show that retraining (abbreviated as RT henceforth) can improve accuracy in the presence of label noise, we propose to apply our proposals in Section 1, viz., full RT and more importantly, consensus-based RT to improve label DP training (because it involves noisy labels). Note that this can be done on top of any label DP mechanism and that too at no additional privacy cost (both the predicted labels and originally provided noisy labels are private). We empirically evaluate full and consensus-based RT on three classification datasets (available on TensorFlow) trained with label DP. These include two vision datasets, namely CIFAR-10 and CIFAR-100, and one language dataset, namely AG News Subset (Zhang et al.,, 2015). All the empirical results are averaged over three different runs. We only provide important experimental details here; the other details can be found in Appendix H.
CIFAR-10/100. We train a ResNet-18 model on CIFAR-10 and CIFAR-100 with label DP. Label DP training is done with the prior-based method of Ghazi et al., (2021) – specifically, Alg. 3 with two stages. Our training set consists of 45k examples and we assume access to a validation set with clean labels consisting of 5k examples which we use for deciding when to stop training, setting hyper-parameters, etc. [^11] In practice, we do not need full access to the validation set. Instead, the validation set can be stored by a secure agent which returns us a private version of the validation accuracy and this will not be too far off from the true validation accuracy when the validation set is large enough. For CIFAR-10 and CIFAR-100 with three different values of $\epsilon$, we list the test accuracies of the baseline (i.e., the method of Ghazi et al., (2021)), full RT and consensus-based RT in Tables 1 and 2, respectively. Notice that consensus-based RT is the clear winner. Also, for the three values of $\epsilon$ in Table 1 (CIFAR-10), the size of the consensus set (used in consensus-based RT) is $\sim 31\%$, $55\%$ and $76\%$, respectively, of the entire training set. The corresponding numbers for Table 2 (CIFAR-100) are $\sim 11\%$, $34\%$ and $56\%$, respectively. So for small $\epsilon$ (high label noise), consensus-based RT comprehensively outperforms full RT and baseline with a small fraction of the training set. Further, in Tables 3 and 4, we list the accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted (= given) labels over the consensus set for CIFAR-10 and CIFAR-100. Please see the table captions for detailed discussion but to summarize, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than the full dataset.
@@
Table 3: CIFAR-10. Accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted (= given) labels over the consensus set. Note that the accuracy of predicted labels over the consensus set $\gg$ accuracy of predicted labels over the entire dataset $\gg$ accuracy of given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than the full dataset ($\sim$ $31\%$, $55\%$ and $76\%$ of the full dataset for $\epsilon=1,2$ and $3$, respectively).
@@
Table 4: CIFAR-100. Accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted (= given) labels over the consensus set. Here the accuracy of predicted labels over the consensus set $\gg$ accuracy of predicted labels over the entire dataset $\approx$ accuracy of given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is much smaller than the full dataset ($\sim$ $11\%$, $34\%$ and $56\%$ of the full dataset for $\epsilon=3,4$ and $5$, respectively). But unlike CIFAR-10 (Table 3), here the accuracy of predicted labels over the entire dataset is not too much better than the given labels. Thus, full RT is only somewhat better than the baseline for CIFAR-100 (see Table 2).
@@
AG News Subset (https://www.tensorflow.org/datasets/catalog/ag_news_subset). This is a news article classification dataset consisting of 4 categories – world, sports, business or sci/tech. We reserve 10% of the given training set for validation and we use the rest for training with label DP. Just like the CIFAR experiments, we assume that the validation set comes with clean labels. We use the small BERT model available in TensorFlow and the BERT English uncased preprocessor; links to both of these are in Appendix H. We pool the output of the BERT encoder, add a dropout layer with probability = 0.2, followed by a softmax layer. We fine-tune the full model. Here, label DP training is done with randomized response. We list the test accuracies of the baseline (i.e., randomized response), full RT and consensus-based RT in Table 5 for three different values of $\epsilon$. Even here, consensus-based RT is the clear winner. For the three values of $\epsilon$ in Table 5, the size of the consensus set (used in consensus-based RT) is $\sim 28\%$, $32\%$ and $38\%$, respectively, of the entire training set. So here, consensus-based RT appreciably outperforms full RT and baseline with less than two-fifths of the entire training set. Finally, in Table 6, we list the accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted (= given) labels over the consensus set. Even here, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This explains why consensus-based RT performs the best, even though the consensus set is much smaller than the full dataset.
@@
Table 6: AG News Subset. Accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted (= given) labels over the consensus set. Conclusions are the same as Table 3.
@@
So in summary, consensus-based RT significantly improves model accuracy. In Appendix I, we show that consensus-based RT also outperforms retraining on samples for which the model is the most confident; this is similar to self-training’s method of sample selection in the semi-supervised setting. Furthermore, going beyond label DP, we show that consensus-based RT is also beneficial in the presence of human annotation errors which can be thought of as “real” label noise in Appendix J. Also, the test accuracies without label DP for all the above datasets are listed in Appendix H.
6 Conclusion and Limitations
In this work, we provided the first theoretical result showing retraining with hard labels can provably increase model accuracy in the presence of label noise. We also showed the efficacy of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) in improving label DP training at no extra privacy cost.
We will conclude by discussing some limitations of our work which also pave the way for some future directions of work. Our retraining analysis is done on an unlabeled dataset different from the initial labeled dataset used for vanilla training (due to certain technical complications). In the future, we would like to analyze retraining on the same labeled dataset. We did not investigate the tightness of the results in Theorem 5 and Remark 4; investigating this is left for future work. Moreover, we have only analyzed full retraining in this work. Given that consensus-based retraining worked very well empirically, we would like to analyze it theoretically later. Another potential extension of our work is to analyze retraining under non-uniform label noise models. Our experiments in this paper are not on very large-scale models or datasets. We hope to test our ideas on larger scale problems in the future.
@@ Appendix A Problem Setting of Figure 1
The setting is exactly the same as the problem setting in Section 4 with $\bm{\theta}^{*}={\mathbf{e}}_{1}$, =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$, $\sigma=1$, $d=50$ and $p=0.4$, except for the distribution of ${\mathrm{a}}^{(1)}$ which is equal to ${\mathrm{x}}^{(1)}$ as $\bm{\theta}^{*}={\mathbf{e}}_{1}$. Specifically for some $\Delta>0$, ${\mathrm{x}}^{(1)}$ is drawn from $\text{Unif}[\Delta,5\Delta]$ with probability $\frac{1}{2}$ or from $\text{Unif}[-5\Delta,-\Delta]$ otherwise ($\text{Unif}[a,b]$ denotes the uniform distribution over the interval $[a,b]$). Note that this distribution is similar in principle to the one assumed in Section 4 (where we did not assume ${\mathrm{a}}^{(1)}$ to have any specific distribution). In Figure 1(a), $\Delta=0.5$ (large separation) and in Figure 1(b), $\Delta=0.3$ (small separation). The number of training samples in each case is 300 and the retraining is done on the same training set on which the model is initially trained. The learned classifiers from vanilla training and retraining are the same as in Section 4 (i.e., eq. 2 and eq. 3, respectively). Finally, the test accuracy of both vanilla training and retraining in Figure 1(b) is $68\%$.
Appendix B Proof of Theorem 1
Proof.
Recall the problem setting at the beginning of Section 4. Without loss of generality, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ (i.e., the canonical basis). In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$. Let ${\mathrm{y}}=\text{sign}({\mathrm{x}}^{(1)})$ be the ground truth label of ${\mathbf{x}}$. Note that proving Theorem 1 is equivalent to proving the following statement. For a sample ${\mathbf{x}}\notin\mathcal{X}_{L}$, the following holds with a probability of at least $\alpha({\mathbf{x}})=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ and at most $\widehat{\alpha}({\mathbf{x}})=1-\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}$:
$\begin{cases}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\geq 0\text{ when }{\mathrm{y}}=1\text{ i.e., }{\mathrm{x}}^{(1)}\geq 0,\\ \langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle<0\text{ when }{\mathrm{y}}=-1\text{ i.e., }{\mathrm{x}}^{(1)}<0.\end{cases}$ (7)
(a) Lower bound $\alpha({\mathbf{x}})$. First note that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$ in our choice of basis for the proof. We will prove the result for the case of ${\mathrm{y}}=1$ or equivalently ${\mathrm{x}}^{(1)}\geq 0$; the result for the other case (i.e., ${\mathrm{y}}=-1$ or equivalently ${\mathrm{x}}^{(1)}<0$) can be derived similarly.
We have:
$\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle=\frac{1}{\sigma^{2}}\Bigg{\{}\frac{1}{n}\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle\Bigg{\}}.$ (8)
Since $\frac{1}{n\sigma^{2}}$ is a constant independent of the training set $\mathcal{T}$, we will analyze
$\varepsilon=\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle.$ (9)
Using the Chernoff bound, we have for any $r<0$:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $=\mathbb{P}\big{(}\exp\big{(}r\varepsilon\big{)}\geq 1\big{)}\leq\mathbb{E}\Big{[}\exp\big{(}r\varepsilon\big{)}\Big{]}.$ (10)
For $k\in[n]$, let $\varepsilon_{k}:=\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle$; thus, $\varepsilon=\sum_{k\in[n]}\varepsilon_{k}$. For any $r\in\mathbb{R}$, we have:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon\big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}r\sum_{k\in[n]}\varepsilon_{k}\big{)}\Big{]}=\mathbb{E}\Big{[}\prod_{k\in[n]}\exp\big{(}r\varepsilon_{k}\big{)}\Big{]}=\Big{(}\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}\big{)}\Big{]}\Big{)}^{n},$ (11)
where the last step follows because of the i.i.d. nature of the data and label noise. Now:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}\big{)}\Big{]}$ $=\mathbb{E}\Big{[}\exp\big{(}r\widehat{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}.$ (12)
Using the definition of $\widehat{\mathrm{y}}_{1}$ from eq. 1, we have:
$\mathbb{E}\Big{[}\exp\big{(}r\widehat{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $=\Bigg{(}(1-p)\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}+p\mathbb{E}\Big{[}\exp\big{(}-r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}\Bigg{)}.$ (13)
So now we need to evaluate $\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ and $\mathbb{E}\Big{[}\exp\big{(}-r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$. Recalling that ${\mathrm{y}}_{1}=\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}$, we have:
$\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}=\mathbb{E}\Bigg{[}\exp\Bigg{(}r\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}\Bigg{(}{\mathrm{x}}_{1}^{(1)}{\mathrm{x}}^{(1)}+\sum_{j=2}^{d}{\mathrm{x}}_{1}^{(j)}{\mathrm{x}}^{(j)}\Bigg{)}\Bigg{)}\Bigg{]},$ (14)
Simplifying the above by using the independence of ${\mathrm{x}}_{1}^{(j)}$’s and recalling that each ${\mathrm{x}}_{1}^{(j)}\sim\mathcal{N}(0,\sigma^{2})$ for $j\in\{2,\ldots,d\}$, we get:
$\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $=\mathbb{E}_{{\mathrm{x}}_{1}^{(1)},{\mathrm{x}}_{1}^{(2)},\ldots,{\mathrm{x}}_{1}^{(d)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\prod_{j=2}^{d}\exp\Big{(}r{\mathrm{x}}^{(j)}\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}{\mathrm{x}}_{1}^{(j)}\Big{)}\Bigg{]}$ (15)
$=\mathbb{E}_{{\mathrm{x}}_{1}^{(1)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\prod_{j=2}^{d}\mathbb{E}_{{\mathrm{x}}_{1}^{(j)}}\Big{[}\exp\Big{(}r{\mathrm{x}}^{(j)}\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}{\mathrm{x}}_{1}^{(j)}\Big{)}\Big{]}\Bigg{]}$ (16)
$=\mathbb{E}_{{\mathrm{x}}_{1}^{(1)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\prod_{j=2}^{d}\exp\Bigg{(}\frac{\big{(}\sigma r{\mathrm{x}}^{(j)}\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}\big{)}^{2}}{2}\Bigg{)}\Bigg{]}$ (17)
$={\mathbb{E}_{{\mathrm{x}}_{1}^{(1)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\Bigg{]}}\exp\Bigg{(}\frac{(\sigma r)^{2}}{2}\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}\Bigg{)}$ (18)
$\leq\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r)^{2}}{2}({\mathrm{x}}^{(1)})^{2}\Bigg{)}\exp\Bigg{(}\frac{(\sigma r)^{2}}{2}\Big{(}\|{\mathbf{x}}\|^{2}-({\mathrm{x}}^{(1)})^{2}\Big{)}\Bigg{)}$ (19)
$=\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}.$ (20)
Equation 17 follows from the standard formula of the moment-generating function of a zero-mean Gaussian. Equation 19 follows because $|{\mathrm{x}}_{1}^{(1)}|$ is SG($\sigma^{2}$) with $\mathbb{E}\big{[}|{\mathrm{x}}^{(1)}|\big{]}=\mu\sigma$ and $\sum_{j=1}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}=\|{\mathbf{x}}\|^{2}$. Since the derivation above holds for any $r\in\mathbb{R}$, we also have:

### v2

@@ 5 Improving Label DP Training with Retraining (RT)
Motivated by our theoretical results in Section 4 which show that retraining (abbreviated as RT henceforth) can improve accuracy in the presence of label noise, we propose to apply our proposals in Section 1, viz., full RT and more importantly, consensus-based RT to improve label DP training (because it involves noisy labels). Note that this can be done on top of any label DP mechanism and that too at no additional privacy cost (both the predicted labels and originally provided noisy labels are private). We empirically evaluate full and consensus-based RT on three classification datasets (available on TensorFlow) trained with label DP. These include two vision datasets, namely CIFAR-10 and CIFAR-100, and one language dataset, namely AG News Subset (Zhang et al.,, 2015). All the empirical results are averaged over three different runs. We only provide important experimental details here; the other details can be found in Appendix G.
CIFAR-10/100. We train a ResNet-18 model on CIFAR-10 and CIFAR-100 with label DP. Label DP training is done with the prior-based method of Ghazi et al., (2021) – specifically, Alg. 3 with two stages. Our training set consists of 45k examples and we assume access to a validation set with clean labels consisting of 5k examples which we use for deciding when to stop training, setting hyper-parameters, etc. [^4] In practice, we do not need full access to the validation set. Instead, the validation set can be stored by a secure agent which returns us a private version of the validation accuracy and this will not be too far off from the true validation accuracy when the validation set is large enough. For CIFAR-10 and CIFAR-100 with three different values of $\epsilon$, we list the test accuracies of the baseline (i.e., the method of Ghazi et al., (2021)), full RT and consensus-based RT in Tables 1 and 2, respectively. Notice that consensus-based RT is the clear winner. Also, for the three values of $\epsilon$ in Table 1 (CIFAR-10), the size of the consensus set (used in consensus-based RT) is $\sim 31\%$, $55\%$ and $76\%$, respectively, of the entire training set. The corresponding numbers for Table 2 (CIFAR-100) are $\sim 11\%$, $34\%$ and $56\%$, respectively. So for small $\epsilon$ (high label noise), consensus-based RT comprehensively outperforms full RT and baseline with a small fraction of the training set. Further, in Table 3, we list the accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted labels (which are the same as the given labels) over the consensus set for CIFAR-10 and CIFAR-100. To summarize, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than full dataset.
@@
Table 3: CIFAR-10 (Top) and CIFAR-100 (Bottom). Accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted labels over the consensus set. Note that the accuracy over the consensus set $\gg$ accuracy of over the entire dataset (with both predicted and given labels). This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than the full dataset ($\sim$ $31\%$, $55\%$ and $76\%$ of the full dataset for $\epsilon=1,2$ and $3$ in the case of CIFAR-10, and $\sim$ $11\%$, $34\%$ and $56\%$ of the full dataset for $\epsilon=3,4$ and $5$ in the case of CIFAR-100).
CIFAR-10 $\epsilon$ Acc. of predicted labels on full dataset Acc. of given labels on full dataset Acc. of predicted labels on consensus set
@@
CIFAR-100 $\epsilon$ Acc. of predicted labels on full dataset Acc. of given labels on full dataset Acc. of predicted labels on consensus set
@@
One may wonder how more model parameters affect the results due to potential overfitting. So here we consider training ResNet-34 (which has more parameters than ResNet-18) on CIFAR-100 with label DP; the setup is exactly the same as our previous experiments with ResNet-18. Please see the results and discussion in Table 4. In summary, consensus-based RT is pretty effective even with ResNet-34 but the amount of improvement here is less compared to ResNet-18 due to overfitting. Additionally, one may wonder if retraining is useful when a label noise-robust technique is used during initial training, i.e., in the baseline (the method of Ghazi et al., (2021)). To that end, we once again train ResNet-34 on CIFAR-100 but with the popular noise-correcting technique of “forward correction” (Patrini et al.,, 2017) applied to the first stage of our baseline (it is not clear how to apply it to the second stage). [^5] Forward correction worked the best among other popular noise-robust techniques such as backward correction (Patrini et al.,, 2017) and noise-robust loss functions (Wang et al.,, 2019; Zhang and Sabuncu,, 2018). Please see the result and discussion for $\epsilon=5$ in Table 5. In summary, retraining can offer gains even after noise-robust training.
Table 4: CIFAR-100 w/ ResNet-34. Test set accuracies (mean $\pm$ standard deviation). Just like Table 2 (ResNet-18), consensus-based RT is the clear winner here. However, note that the performance and amount of improvement with ResNet-34 is worse than ResNet-18 due to more overfitting because of more parameters; this is expected.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $17.53\pm 0.05$ $18.20\pm 0.29$ $\textbf{21.63}\pm 0.31$
4 $37.53\pm 1.58$ $39.33\pm 1.37$ $\textbf{43.87}\pm 1.62$
5 $51.13\pm 0.69$ $52.33\pm 0.38$ $\textbf{55.43}\pm 0.42$
Table 5: Noise-robust “forward correction” (Patrini et al.,, 2017) applied to baseline (CIFAR-100 w/ ResNet-34). Test set accuracies (mean $\pm$ standard deviation). We see that retraining can yield improvement even after noise-robust training, although the amount of improvement is less compared to Table 4 (no noise-correction).
$\epsilon$ Baseline Full RT Consensus-based RT
5 $53.83\pm 0.83$ $54.43\pm 0.48$ $\textbf{56.60}\pm 0.43$
AG News Subset (https://www.tensorflow.org/datasets/catalog/ag_news_subset). This is a news article classification dataset consisting of 4 categories – world, sports, business or sci/tech. We reserve 10% of the given training set for validation and we use the rest for training with label DP. Just like the CIFAR experiments, we assume that the validation set comes with clean labels. We use the small BERT model available in TensorFlow and the BERT English uncased preprocessor; links to both of these are in Appendix G. We pool the output of the BERT encoder, add a dropout layer with probability = 0.2, followed by a softmax layer. We fine-tune the full model. Here, label DP training is done with randomized response. We list the test accuracies of the baseline (i.e., randomized response), full RT and consensus-based RT in Table 6 for three different values of $\epsilon$. Even here, consensus-based RT is the clear winner. For the three values of $\epsilon$ in Table 6, the size of the consensus set (used in consensus-based RT) is $\sim 28\%$, $32\%$ and $38\%$, respectively, of the entire training set. So here, consensus-based RT appreciably outperforms full RT and baseline with less than two-fifths of the entire training set. Finally, in Table 7, we list the accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted labels (= given labels) over the consensus set. Even here, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This explains why consensus-based RT performs the best, even though the consensus set is much smaller than the full dataset.
@@
Table 7: AG News Subset. Accuracies over the entire dataset and over the consensus set. Conclusions are the same as Table 3.
@@
So in summary, consensus-based RT significantly improves model accuracy.
Additional Empirical Results in the Appendix. In Appendix H, we show that consensus-based RT also outperforms retraining on samples for which the model is the most confident; this is similar to self-training’s method of sample selection in the semi-supervised setting. In Appendix I, we show that RT is beneficial even without a clean validation set. Furthermore, going beyond label DP, we show that consensus-based RT is also beneficial in the presence of human annotation errors which can be thought of as “real” label noise in Appendix J.
6 Conclusion
In this work, we provided the first theoretical result showing retraining with hard labels can provably increase model accuracy in the presence of label noise. We also showed the efficacy of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) in improving label DP training at no extra privacy cost. We will conclude by discussing some future directions of work. Our theoretical results in this work focus on the full retraining scheme. Given that consensus-based retraining worked very well empirically, we would like to analyze it theoretically as a future work. Another potential extension of our work is to analyze retraining under non-uniform label noise models. We also hope to test our ideas on larger scale models and datasets in the future.
@@ Appendix A Problem Setting of Figure 1
The setting is exactly the same as the problem setting in Section 4 with ${\bm{{\mu}}}=\gamma\bm{e}_{1}$, $\bm{\Sigma}=\bm{I}-\bm{e}_{1}\bm{e}_{1}^{\sf T}$, $d=50$ and $p=0.4$. In the direction of ${\bm{{\mu}}}$, we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}=y(1+u)\gamma^{2}$. We consider balanced classes (i.e., $\pi_{+}=\pi_{-}=\frac{1}{2}$) and choose $u\sim{\sf Unif}[0,4]$, the uniform distribution over the interval $[0,4]$. In Figure 1(a), $\gamma^{2}=0.5$ (large separation) and in Figure 1(b), $\gamma^{2}=0.3$ (small separation). The number of training samples in each case is 300 and the retraining is done on the same training set on which the model is initially trained. The learned classifiers from vanilla training and retraining are the same as in Section 4 (i.e., eq. 4.3 and eq. 4.7, respectively). Finally, the test accuracy of both vanilla training and retraining in Figure 1(b) is $68\%$.
Appendix B Proof of Theorem 4.1
We first prove the upper bound $\alpha_{0}(\bm{x})$ and then the lower bound $\widetilde{\alpha}_{0}(\bm{x})$.
B.1 Upper bound $\alpha_{0}(\mathbf{x})$
Let $\xi_{i}=y_{i}\widehat{y}_{i}$. So $\xi_{i}=1$ with probability $1-p$ and $\xi_{i}=-1$ with probability $p$. Under our data model, for each sample we can write $\bm{x}_{i}=y_{i}(1+u_{i}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\bm{z}_{i}$ with $\bm{z}_{i}\sim{\mathcal{N}}(0,\bm{I}_{d})$. Hence,
$n{\widehat{\bm{\theta}}}_{0}=\sum_{i\in[n]}\widehat{y}_{i}\bm{x}_{i}=(\sum_{i\in[n]}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\sum_{i\in[n]}\widehat{y}_{i}\bm{\Sigma}^{1/2}\bm{z}_{i}\,.$
We also note that
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)$ $=\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle\leq 0)=\mathbb{P}\left((\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle+\langle\bm{x},\sum_{i\in[n]}y\widehat{y}_{i}\bm{\Sigma}^{1/2}{\bm{z}}_{i}\rangle\leq 0\right)$
Define $\tilde{\bm{z}}:=\sum_{i\in[n]}y\widehat{y}_{i}\bm{\Sigma}^{1/2}{\bm{z}}_{i}$. Conditioning on $\{\widehat{y}_{i}\}_{i\in[n]}$ we have $\tilde{\bm{z}}\sim{\mathcal{N}}(\bm{0},n\bm{\Sigma})$, and so $\langle\bm{x},\tilde{\bm{z}}\rangle\sim{\mathcal{N}}(\bm{0},n\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})$.
We write